OpenStudy (anonymous):

I am trying to figure out what number would make this problem a perfect squared trimomial? x^2+13x

OpenStudy (campbell_st):

well a perfect square (x + a)^2 = x^2 + 2ax + a^2 so use the coefficient of the middle term, which is 13... divide it by 2 and then square the answer. hope it helps

OpenStudy (anonymous):

I guess I am just not understanding what my teacher is asking for because 13 can not be squared

OpenStudy (anonymous):

the unit I am in is quadratic equations ugg I am frustrated

OpenStudy (campbell_st):

if you look at the perfect square (x + a)^2 = x^2 + 2ax + a^2 13 isn't a square number... I think is what you were trying to say. the task is to find a so you know the middle term is 13x so 2ax = 13x when comparing the middle terms that means that a = 13/2 now this is a number that can be squared \[(\frac{13}{2})^2 = \frac{13^2}{2^2}\] and that's what you add to make a perfect square

OpenStudy (campbell_st):

well I'd expect its \[x^2 + 13x + c\] so the constant is half the middle coefficient squared... as shown above

OpenStudy (anonymous):

The question reads as " Determine the constant that should be added to the binomial so that it becomes a perfect squared trinomial.' \[X^2+13x\]

OpenStudy (anonymous):

ok I think i get it

OpenStudy (campbell_st):

so if you have 13/2 = a then (x + a)^2 = x^2 + 2ax + a^2 so (x + 13/2)^2 = x^2 + 2(13/2)x + (13/2)^2

OpenStudy (campbell_st):

and 2(13/2)x = 13x so its x^2 + 13x + (13/2)^2

OpenStudy (anonymous):

YES!!! THANK YOU! Now i get it